. An infectious disease is introduced into a school with 1200 students. Fortunately, the disease only lasts a few days, and all of the infected students recover and then have long lasting immunity against the disease (they can only be infected once). The incidence rate for the outbreak, which begins the first week of school, remains constant throughout the school year at 1 new case per hundred uninfected students per week. By the end of the forty-week school year, how many students would you expect to have contracted the disease? Use the formula for approximating the cumulative incidence from the incidence rate.
Answer #1

The formula for cumulative incidence is

CI = IR × T

Where CI is cumulative incidence

IR is incidence rate

And T is time

This formula is used mainly when rate is low and constant and many person or most of the population is at risk

So,

in this case is 1/100/weeks × 40weeks = 40 per 100 student

So there are 1200 student in the school so

40 × 12 = 480

So 480 is the correct answer to this question

By the end of 40 weeks 480 student would have been in contact the disease

Hope you find this helpful thankyou ☺️?

So